Ellen who is 20 years old has a grandmother who is 65. How many yeas ago was the grandmother 6 times as old as Ellen.
Answer by ankor@dixie-net.com(22740)

Ellen who is 20 years old has a grandmother who is 65.
How many yeas ago was the grandmother 6 times as old as Ellen?
:
let y = no. of yrs ago this was true
:
65 - y = 6(20 - y)
65 - y = 120 - 6y
-y + 6y = 120 - 65
5y = 55
y = 55/5
y = 11 yrs ago
:
:
You can check this
11 yrs ago, Grandma was 54 and Ellen was 9; 9*6 = 54